OpenStudy (i_ask_alot):

What is the circumference of a 12-inch diameter pizza? Leave the answer in terms of π.

OpenStudy (i_ask_alot):

@TheSmartOne

OpenStudy (michele_laino):

hint: we can apply this formula: \[\huge C = 2\pi r = \pi d\] where \(d\) is the diameter

OpenStudy (i_ask_alot):

im confused

OpenStudy (michele_laino):

why?

OpenStudy (i_ask_alot):

what does r represent?

OpenStudy (michele_laino):

we have \(2r=d\) so we can write: \[\huge C = 2\pi r = \left( {2r} \right)\pi = d\pi = \pi d\] r is the radius of the circumference

OpenStudy (i_ask_alot):

is it 12π

OpenStudy (michele_laino):

that's right!

OpenStudy (i_ask_alot):

Thanks

OpenStudy (michele_laino):

:)